Can somebody PLEASE help me with this question, I have tryed to do it so many times, but just can't seem to figure it out. Any help would be appreciated!:)
a.)For the sequence 0,3,8,15,24....what would be the next three terms?
b.)What would be the function f(x) that would correspond to this sequence?
c.)what would the genera term tn for the nth term of this sequence be?
d.)How would you describe this sequence recursively, and write a recursive formula for this sequence?
e.)How would you find the 13th term of this sequence, using your formula from b.?

Answer by stanbon(57347)

Can somebody PLEASE help me with this question, I have tryed to do it so many times, but just can't seem to figure it out. Any help would be appreciated!:)
a.)For the sequence 0,3,8,15,24....what would be the next three terms?
Notice that each number is 1 less than a perfect square.
---
1st term: 1^2-1
2nd term: 2^2-1
3rd term: 3^2-1
....
6th term: 6^2-1 = 35
7th term: 6^2-1 = 48
8th term: 8^2-1 = 63
-------------------------------------------
b.)What would be the function f(x) that would correspond to this sequence?
f(x) = x^2-1
--------------------
c.)what would the general term tn for the nth term of this sequence be?
t(n) = n^2-1
-------------------------------
d.)How would you describe this sequence recursively, and write a recursive formula for this sequence?
a(n) = a(n-1)+2n-1
----------------------------
e.)How would you find the 13th term of this sequence, using your formula from b.?
f(13) = 13^2-1 = 168
============================
Cheers,
Stan H.